# FedMLLM: Federated Fine-tuning MLLM on Multimodal Heterogeneity Data

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have made significant advancements, demonstrating powerful capabilities in processing and understanding multimodal data. Fine-tuning MLLMs with Federated Learning (FL) allows for expanding the training data scope by including private data sources, thereby enhancing their practical applicability in privacy-sensitive domains. However, current research remains in the early stage, particularly in addressing the multimodal heterogeneities in real-world applications. In this paper, we introduce a benchmark for evaluating various downstream tasks in the federated fine-tuning of MLLMs within multimodal heterogeneous scenarios, laying the groundwork for the research in the field. Our benchmark encompasses two datasets, five comparison baselines, and four multimodal scenarios, incorporating over ten types of modal heterogeneities. To address the challenges posed by modal heterogeneity, we develop a general FedMLLM framework that integrates four representative FL methods alongside two modality-agnostic strategies. Extensive experimental results show that our proposed FL paradigm improves the performance of MLLMs by broadening the range of training data and mitigating multimodal heterogeneity. Code is available at https://github.com/1xbq1/FedMLLM
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), which regard Large Language Models (LLMs) as core components for performing various multimodal tasks, have rapidly emerged as a research hotspot [39]. Their exceptional performance and potential as a pathway to artificial general intelligence contribute to this trend. However, the high computational costs and the requirement for large amounts of training data are significant challenges that hinder the further development and application of MLLMs in the real world [37, 24]. For the former, various lightweight MLLMs [37, 23] and Parameter Efficient Fine-Tuning (PEFT) [11, 20] methods, such as Low-Rank Adaptation (LoRA) [11], have been investigated. For the latter, a practical solution is to extend the data scope from public datasets to unused private data via Federated Learning (FL) [27, 15].
However, there is limited research on fine-tuning MLLMs under the FL framework [43], with most efforts focused on integrating LLMs with FL [38, 19, 42]. Compared to LLMsFL, MLLMsFL supports more modal inputs, making it more applicable to real-world scenarios [39]. For example, in natural disasters or emergency situations (e.g., earthquakes, floods), it is essential to integrate multi-modal data such as images or videos (on-site photos), text (social media reports), and audio (distress signals). MLLMs+FL enable distributed collaboration across organizations while safeguarding data privacy, offering a more comprehensive understanding of disasters through multi-modal modeling. In contrast, LLMs+FL are limited to text, missing key information from images and audio. However, the multimodal data, as the input of FL framework of multi-client collaborative training, also brings about more complex modal heterogeneity. As shown in Fig. 1, multimodal scenarios can be categorized as the aligned modality [9], and the non-aligned modality (i.e., missing modality [26, 3], cross modality [36], and hybrid modality [40, 5]). Only the heterogeneity of the aligned modality is consistent with that of a single modality, while the non-aligned multimodal situations exhibit additional modal heterogeneity.
To eliminate the bias caused by modal heterogeneity, we design two simple yet effective modal-agnostic strategies based on existing FL methods, as illustrated in Fig. 2. First, the prompt strategy instructs LLMs to perform tasks without considering modality, e.g., “…, without considering the modality.”, which considers the effect of instructions on LLMs at the data level [33]. Second, the regularization strategy constrains the update direction. As the client moves from aligned to missing modalities, eventually to extreme single-modality cases, the risk of overfitting to a specific modality gradually increases. To address this, the adaptive regularization strategy adjusts consistency with global weights based on the severity of missing modalities, with stronger consistency for more missing modalities. Meanwhile, only the middle layers are regularized, preserving the ability of the low and high layers to extract modal features and provide task-specific representations.
Combined with the above modal-agnostic strategies, we comprehensively explore the pipeline of federated fine-tuning of MLLM on decentralized multimodal downstream data. Unlike [43], which applies MLLM solely to assist FL in alleviating traditional data heterogeneity and long-tail distribution problems, we focus more on addressing the issue of diverse modal heterogeneity in multimodal scenarios that are closely related to MLLMs. As shown in Fig. 2, the proposed pipeline supports four multimodal scenarios: i. Aligned Modal Scenario refers to all samples of each client existing in an aligned multimodal form. ii. Missing Modal Scenario occurs when a certain modality is absent in some samples of each client. iii. Cross Modal Scenario involves each client having only one modality, while all clients together cover multiple modalities. iv. Hybrid Modal Scenario refers to some clients containing only one modality, while others contain a combination of multiple modalities [9]. In total, we construct over ten cases of modal heterogeneity across two datasets, based on varying heterogeneity degrees and the proportions of non-aligned multi-modalities. Finally, the application of FL in MLLMs significantly improves performance compared to local training, as shown in the four modal results of Fig. 2. Additionally, the two modality-agnostic strategies proposed for the modal heterogeneity problem also contribute positively.
@@
Overall, our contributions can be summarized in three folds. 1) We explore the fine-tuning MLLM benchmark on decentralized multimodal data to better adapt to various downstream tasks in comprehensive multimodal federated scenarios, which highlights a potential development direction for MLLMs. 2) We build a general FedMLLM framework to better address the challenges of multimodal heterogeneity by elegantly incorporating two modality-agnostic strategies on top of multiple FL baselines. 3) We conduct extensive experiments, related to 2 downstream datasets, 5 comparison baselines, and 4 multimodal scenarios, to show that FL paradigm enhances MLLM performance by broadening the training data scope without bells and whistles.
2 Related Work
2.1 Multimodal Large Language Models
The rapid development of Large Language Models (LLMs) has significantly driven the research surge in Multimodal Large Language Models (MLLMs) [39]. Notably, the emergence of commercial models such as GPT-4 [28] and Gemini [32] has attracted widespread attention. Simultaneously, increasingly powerful open-source models [21, 7, 44] have also emerged. BLIP-2 [21] is well-known for its Q-former module, with promising VQA results. Building on BLIP-2 and Vicuna [7], MiniGPT-4 [44], derived from GPT-4 [28], is proposed for image-text pairs. Unlike previous methods, LLaVA [24] is trained on multimodal instruction data generated by GPT-4 to enhance its zero-shot capability.
2.2 Federated Learning
As a classic framework for privacy protection, Federated Learning (FL) typically allows multiple parties to co-train while keeping the data local [12]. The original FedAvg [27] method aggregates model weights from different clients solely based on sample proportions, which makes it challenging to effectively address complex data heterogeneity. As a result, various variants have emerged: some constrain the update direction of the local model on the client side [22, 16], while others optimize the aggregation method on the server side [30, 10, 31].
2.3 Multimodal Federated Learning
Unlike the more established single-modal FL, Multimodal Federated Learning (MFL) has significant potential for exploration [9]. Existing MFL methods can also be categorized based on the different modal scenarios in our benchmark. In aligned modal scenarios, MFL techniques resemble those of single-modal FL, focusing on client-specific and shared knowledge [4]. Some methods also draw from multimodal learning, especially in feature fusion [9]. In missing modal scenarios, some MFL methods rely on pre-trained generative models to fill the gaps in modality [3]. In cross modal scenarios, some MFL methods employ a two-branch structure to separately constrain the learning of specific and shared modalities [36, 35, 29], while hybrid modal MFL methods utilize various graph techniques to capture complex correlations among clients [5, 6].
@@ 3 Multimodal Scenarios Construction
Aligned Modal Scenario. The dataset is sampled using Dirichlet distribution, similar to the traditional partitioning method [9], with coefficients adjusted to control the distribution of samples among different clients. Specifically, for each label, Dirichlet distribution is used to allocate the number of samples to each client; the smaller the coefficient, the greater the difference in sample sizes between clients, resulting in increased data heterogeneity. Here, each client’s sample contains all modalities.
Missing Modal Scenario. The dataset first assigns samples to different clients according to the partitioning strategy of aligned modal scenarios. Currently, the samples in each client are full-modal. By using a uniform distribution for each modality and setting a missing rate of to determine whether to remove that modality, the situation where the samples in each client have different missing modalities is achieved.
Cross Modal Scenario. The dataset is first allocated according to a partitioning strategy similar to the aligned modal scenario and then divided based on the ratio ({, , }) of image-only to text-only clients. For example, in the ratio, 3 out of 10 clients are randomly assigned the image modality, while the other 7 clients receive the text modality. The other ratios follow the same pattern. This approach allows each client to have a single modality while ensuring that all clients collectively have multiple modalities.
Hybrid Modal Scenario. After applying a partitioning method consistent with the aligned modality scenario, the dataset follows a Bernoulli distribution for each client’s modality retention probability , which is the same for each modality. Once the partitioning is complete, a single client may consist of either a single modality or a combination of multiple modalities, with internal samples maintaining consistency in modality. To avoid an invalid dataset missing both modalities on the client side, in such cases, either the image or text modality is randomly retained with a 50% probability.
As shown in Fig. 3 and Fig. 4, we present visualizations of heterogeneity in the aligned modal scenario on the CrisisMMD dataset and of multimodal heterogeneities across four multimodal scenarios on the Hateful-Memes dataset, respectively. More visualization results can be found in the Supplementary Material. Overall, these four multimodal scenarios initially simulate simple situations with various modal inputs in real life.
@@ 4.1 Overview of FedMLLM
4.2 Modality-agnostic Strategies
Prompt Strategy. In decentralized multimodal data scenarios, different modality conditions among clients further complicate the data heterogeneity problem. Lacking modalities may introduce additional modality bias into the model’s learning process, affecting the convergence of the global model. A common solution is to address this issue at the data source by using a large-scale pre-trained model to supplement the modality [3]. Nevertheless, it often incurs additional costs. To avoid these costs, we aim to mitigate the impact of modality changes on task outcomes by explicitly prompting model training. Inspired by this, we augment the original prompt by introducing a modality-agnostic prompt strategy to enhance the model’s robustness and generalization ability, as shown in Tab. 1.
Regularization Strategy. The regularization term originally introduced by FedProx [22] constrains the direction of local updates toward the global model, serving as a general method for mitigating heterogeneity. However, the average aggregation strategies used in FedProx and FedAvg [27] can exacerbate instability during federated fine-tuning of MLLM, leading to a gradual increase in loss and its variance (see in Supplementary Material). In contrast, FedAvgM employs server momentum [10], while FedAdam, FedYogi, and FedAdagrad use adaptive optimizers [30], allowing them to handle instability more effectively.
@@ 4.2 Modality-agnostic Strategies
| Local | 71.72 | 55.56 | 67.38 | 45.78 | 67.71 | 37.67 |
| FedYogi | 74.56 | 61.21 | 71.53 | 61.56 | 71.13 | 60.90 |
| FedAdam | 74.67 | 61.04 | 73.17 | 59.52 | 71.74 | 61.01 |
| FedAvgM | 74.39 | 56.69 | 74.55 | 56.76 | 62.51 | 32.09 |
| FedAdagrad | 73.90 | 60.88 | 71.99 | 60.66 | 72.69 | 59.61 |
@@
| Local | 66.83 | 33.34 | 66.94 | 31.88 | 66.90 | 31.56 |
| FedYogi | 69.96 | 60.06 | 69.12 | 59.16 | 74.17 | 54.57 |
| FedAdam | 72.52 | 60.61 | 72.35 | 57.96 | 73.15 | 54.33 |
| FedAvgM | 63.07 | 33.17 | 59.92 | 34.57 | 58.69 | 32.99 |
| FedAdagrad | 72.67 | 56.65 | 71.05 | 56.30 | 70.94 | 56.17 |
@@
| Local | 67.04 | 25.92 | 66.38 | 28.82 | 67.38 | 31.42 |
| FedYogi | 67.02 | 47.46 | 75.09 | 58.03 | 76.02 | 60.97 |
| FedAdam | 67.35 | 51.88 | 74.78 | 58.20 | 74.56 | 60.91 |
| FedAvgM | 61.50 | 31.31 | 63.91 | 38.67 | 58.77 | 33.41 |
| FedAdagrad | 69.33 | 49.22 | 72.47 | 51.21 | 72.35 | 51.55 |
@@
| Local | 67.10 | 35.37 | 67.66 | 33.52 | 67.68 | 33.70 |
| FedYogi | 66.36 | 58.60 | 69.60 | 60.82 | 70.02 | 58.57 |
| FedAdam | 70.43 | 60.84 | 70.68 | 61.91 | 71.26 | 59.39 |
| FedAvgM | 60.43 | 47.92 | 62.12 | 43.27 | 60.35 | 39.07 |
| FedAdagrad | 69.69 | 57.36 | 72.18 | 57.87 | 69.93 | 53.45 |
where is an adaptive coefficient, is the layer-wise mask, is the transpose operation, is a layer-count hyperparameter, and represents the server-side aggregated LoRA weight. In cross modal scenario, suppose is set to . When the missing rate is in missing modal scenario, is linearly adjusted to to adaptively adjust the strength of agreement with the global model. The hybrid modal scenario combines aligned, missing, and cross modal scenarios. Thus, the setting follows these scenarios, with no regularization in the aligned case. In practice, is calculated from the client’s private dataset.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have made significant advancements, demonstrating powerful capabilities in processing and understanding multimodal data. Fine-tuning MLLMs with Federated Learning (FL) allows for expanding the training data scope by including private data sources, thereby enhancing their practical applicability in privacy-sensitive domains. However, current research remains in the early stage, particularly in addressing the multimodal heterogeneities in real-world applications. In this paper, we introduce a benchmark to evaluate the performance of federated fine-tuning of MLLMs across various multimodal heterogeneous scenarios, laying the groundwork for future research in the field. Our benchmark includes two lightweight MLLMs, two downstream tasks, three evaluation metrics, and five datasets across three domains, along with six comparison baselines, covering over ten types of modality heterogeneities across four multimodal scenarios. To address the challenges posed by multimodal heterogeneity, we develop a general FedMLLM framework that integrates classic FL methods alongside two modality-agnostic strategies. Extensive experimental results show that our proposed FL paradigm improves the performance of MLLMs by broadening the range of training data and mitigating multimodal heterogeneity. Code is available in supplementary materials.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), which use Large Language Models (LLMs) as core components for various multimodal tasks, have rapidly emerged as a research hotspot [50]. The exceptional performance of MLLMs and their potential as a pathway to artificial general intelligence have driven this trend. However, the decreasing amount of available public data [42] is an urgent issue that hinders the further development and application of MLLMs in the real world [47, 32]. To expand the available data scope, a practical solution is to fine-tune existing MLLMs, trained on public data, with unused private data via Federated Learning (FL) [36, 18].
However, there is limited research on fine-tuning MLLMs under FL [54, 44]. There are two possible reasons. First, the large parameter size of MLLMs [43] demands significant computational resources for fine-tuning, while FL typically assumes clients have limited resources [26]. Excitingly, recent solutions, such as lightweight MLLMs [47, 28], PEFT methods [13, 24], and model quantization [29], offer support for training MLLMs in FL environments. In this work, we successfully fine-tune MLLMs on the client side of FL, using quantized, lightweight MLLMs [47] and training them via LoRA [13] on a single GPU. This approach is affordable in cross-silo FL settings [14], but cross-device FL [35] remains unfeasible and requires further exploration in future work.
Another possible reason is that current research has not focused on the different challenges faced by MLLMsFL compared to LLMsFL, and thus most studies have concentrated on LLMsFL [48, 22, 53]. Compared to LLMsFL, MLLMsFL supports more modal inputs, making them more applicable to real-world privacy-preserve scenarios [50]. For example, in natural disasters or emergencies (e.g., earthquakes, floods), integrating multimodal data such as images or videos (on-site photos), text (social media reports), and audio (distress signals) is essential for quickly analyzing the situation [2]. Additionally, in hospitals [23], besides text-based medical records, CT images are available for doctors to comprehensively analyze a patient’s condition—an ability that LLMs, which process only text, cannot achieve. Meanwhile, real-world data is influenced by various conditions and situations, making it unrealistic for all clients to have complete modality samples. The absence of certain modalities in samples across clients introduces a new challenge in MLLMsFL, i.e., multimodal heterogeneity.
To comprehensively explore heterogeneity in multimodal data, especially multimodal heterogeneity, we construct four modality scenarios based on existing datasets. In Fig. 1, multimodal data can be categorized into aligned modality [10] and non-aligned modality (i.e., missing modality [34, 4], cross modality [46], and hybrid modality [51, 6]) based on the modality distribution of samples. When federated fine-tuning MLLMs on aligned modality data, client data heterogeneity resembles that in LLM fine-tuning with single-modal text data. However, training MLLMs on non-aligned modality data introduces multimodal heterogeneity. To intuitively demonstrate the heterogeneity in multimodal data, as shown in Figs. 3, 4 and 5, we conduct a visual analysis from multiple perspectives on the multimodal scenarios constructed from different datasets. Among them, visualizations of modality count, types, and feature levels further emphasize the presence of multimodal heterogeneity. More visualizations are in Suppl.
Building on the above multimodal scenarios, we propose a benchmark for federated fine-tuning MLLMs. Specifically, four multimodal scenarios covering more than ten types of modality heterogeneity across five datasets are first constructed. These datasets include classification and medical visual question-answer tasks spanning the domains of social media, natural disasters, and medical knowledge. A general FedMLLM framework based on LoRA is then presented for training and testing on these multimodal scenarios. Using FedMLLM, we compare six baselines and two lightweight MLLMs across three evaluation metrics: AUC, F1, and Accuracy. Additionally, we address the bias introduced by multimodal heterogeneity using two simple yet effective modal-agnostic strategies: improved prompts and regularization techniques. The overview of FedMLLM is as illustrated in Fig. 2.
@@
Overall, our contributions can be summarized in three folds. 1) We explore the fine-tuning MLLM benchmark on decentralized multimodal data to better adapt to downstream tasks in comprehensive multimodal federated scenarios, which highlights a potential development direction for MLLMs. 2) We build a general FedMLLM framework to better address the challenges of multimodal heterogeneity by elegantly incorporating two modality-agnostic strategies on top of multiple FL baselines. 3) We conduct extensive experiments, related to 5 downstream datasets, 6 comparison baselines, and 4 multimodal scenarios, to show that FL paradigm enhances MLLM performance by broadening the training data scope without bells and whistles.
@@ 2 Multimodal Scenarios Construction
Aligned Modal Scenario. The dataset is partitioned using a Dirichlet distribution, similar to the traditional method [10], with coefficients controlling sample distribution among clients. A smaller coefficient increases differences in sample counts across clients. For classification tasks, sampling is based on class labels; for non-classification tasks, it depends on data types, user IDs, etc. Each client’s sample includes all modalities.
Missing Modal Scenario. Building on the Aligned Modal Scenario, a uniform distribution is applied to each modality per sample with a missing rate of to determine modality removal, resulting in samples with different missing modalities across clients. Both the Aligned and Missing Modal Scenarios require the constructed dataset to have complete modalities.
Cross Modal Scenario. This scenario involves two approaches: one for a full-modal dataset (Hateful-Memes, CrisisMMD), where samples are allocated to clients using a Dirichlet distribution, followed by random assignment of image-only and text-only clients in proportion. The other approach is for datasets with only the image modality (VQA-RAD) or text modality (MedAlpaca), where clients are randomly assigned in proportion and the dataset is evenly distributed to the corresponding clients. The proportions used in this study are ({, , }), where I and T denote image-only and text-only clients.
Hybrid Modal Scenario. Each modality per client follows a Bernoulli distribution with retention probabilities . This approach applies to both full-modal and single-modal datasets. After partitioning, clients may have either single-modal or multi-modal data. If both modalities are missing on the client side, one modality is randomly retained with a 50% probability.
@@ 3.1 Overview of FedMLLM
| Dataset | Domain | Evaluate metric | Task | Modality | Training set | Testing set |
| Hateful-Memes [20] | Social media | AUC | Classification | Image&Text | 8,500 | 1,000 |
| CrisisMMD [2] | Natural disasters | F1 | Classification | Image&Text | 13,608 | 2,238 |
| MedAlpaca [11] | Medical knowledge | – | Med-QA | Text | 33,951 | – |
| VQA-RAD [23] | Medical knowledge | Accuracy | Med-VQA | Image | 1,797 | 451 |
| SLAKE [30] | Medical knowledge | Accuracy | Med-VQA | Image | – | 2,094 |
@@
| Zero-shot | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 |
| Local | 70.930.68 | 55.270.29 | 67.100.57 | 47.341.56 | 65.911.27 | 39.121.45 |
| FedYogi | 75.740.88 | 60.951.06 | 72.501.46 | 60.820.58 | 74.320.82 | 61.460.95 |
| FedAdam | 74.021.42 | 61.800.75 | 73.242.43 | 59.120.41 | 72.971.23 | 60.560.56 |
| FedAvgM | 74.081.35 | 58.281.77 | 72.941.61 | 56.870.11 | 64.221.71 | 54.220.42 |
| FedAdagrad | 74.530.55 | 60.770.12 | 73.341.99 | 60.430.23 | 71.800.89 | 60.210.60 |
@@
| Zero-shot | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 |
| Local | 66.720.11 | 33.270.08 | 66.840.10 | 31.990.11 | 66.800.11 | 31.780.22 |
| FedYogi | 76.580.90 | 60.070.01 | 76.750.55 | 56.761.70 | 75.120.70 | 53.780.80 |
| FedAdam | 76.820.05 | 59.541.07 | 77.280.52 | 57.640.32 | 75.581.72 | 54.820.49 |
| FedAvgM | 60.622.45 | 32.640.53 | 57.901.46 | 32.971.61 | 57.341.35 | 31.721.27 |
| FedAdagrad | 71.970.56 | 56.960.31 | 70.640.32 | 57.481.18 | 70.590.27 | 55.550.63 |
3.2 Modality-agnostic Strategies
Prompt Strategy. In decentralized multimodal data scenarios, different modality conditions among clients further complicate the data heterogeneity problem. Lacking modalities may introduce additional modality bias into the model’s learning process, affecting the convergence of the global model. A common solution is to address this issue at the data source by using a large-scale pre-trained model to supplement the modality [4]. Nevertheless, it often incurs additional costs. To avoid these costs, we aim to mitigate the impact of modality changes on task outcomes by explicitly prompting model training. Inspired by this, we augment the original prompt (Original prompt: Is the content hateful based on the text and image?) by introducing a modality-agnostic prompt (Improved prompt: Is the content hateful, without considering the modality?) strategy to enhance the model’s robustness and generalization ability. More detailed prompt templates can be found in Suppl.
Regularization Strategy. The regularization term originally introduced by FedProx [27] constrains the direction of local updates toward the global model, serving as a general method for mitigating heterogeneity. However, the average aggregation strategies used in FedProx and FedAvg [36] can exacerbate instability during federated fine-tuning of MLLM, leading to a gradual increase in loss and its variance (see in Suppl.). In contrast, FedAvgM employs server momentum [12], while FedAdam, FedYogi, and FedAdagrad use adaptive optimizers [39], allowing them to handle instability more effectively.
where is a hyperparameter related to the degree of modality missing, is the layer-wise mask, is the transpose operation, is a layer-count hyperparameter, and represents the server-side aggregated LoRA weight. More details are in Suppl.
@@ 3.2 Modality-agnostic Strategies
| Zero-shot | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 |
| Local | 66.850.20 | 26.360.43 | 66.390.01 | 29.590.77 | 67.570.20 | 31.140.29 |
| FedYogi | 68.711.69 | 49.301.75 | 73.931.16 | 55.532.50 | 75.940.08 | 59.201.77 |
| FedAdam | 65.711.64 | 53.170.96 | 74.440.34 | 56.621.58 | 74.950.39 | 59.221.69 |
| FedAvgM | 58.202.36 | 32.611.30 | 61.572.34 | 38.621.66 | 57.401.37 | 33.460.05 |
| FedAdagrad | 71.261.93 | 48.020.87 | 72.480.05 | 49.970.90 | 71.570.77 | 51.130.42 |
@@
| Zero-shot | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 | 66.570.00 | 24.200.00 |
| Local | 67.050.31 | 35.840.23 | 67.820.16 | 34.000.42 | 67.440.19 | 34.180.20 |
| FedYogi | 68.411.65 | 59.730.96 | 68.550.71 | 63.041.63 | 71.521.13 | 58.960.57 |
| FedAdam | 72.171.62 | 60.020.63 | 70.120.54 | 61.011.12 | 72.360.97 | 59.930.60 |
| FedAvgM | 61.551.33 | 46.261.78 | 60.012.03 | 42.771.05 | 58.931.98 | 39.440.57 |
| FedAdagrad | 68.770.63 | 57.580.32 | 72.960.54 | 58.911.03 | 69.260.47 | 53.840.46 |
| Method | Scenario | VQA-RAD | SLAKE | ||||
| Open | Closed | Overall | Open | Closed | Overall | ||
| Zero-shot | 44.690.00 | 65.070.00 | 56.980.00 | 65.550.00 | 64.110.00 | 64.950.00 | |
| Local | 42.740.09 | 69.970.13 | 59.160.05 | 58.710.68 | 65.830.48 | 61.630.22 | |
| FL | 45.250.56 | 68.021.11 | 58.980.89 | 53.700.35 | 66.150.24 | 58.670.12 | |
| Local | 43.860.43 | 70.040.55 | 59.640.32 | 58.190.20 | 65.190.37 | 60.980.16 | |
| FL | 44.420.84 | 71.140.18 | 60.530.22 | 52.550.23 | 63.040.60 | 56.740.09 | |
| Local | 41.900.29 | 71.690.31 | 59.870.15 | 59.220.33 | 62.320.27 | 60.460.21 | |
| FL | 41.900.75 | 73.530.20 | 60.980.31 | 52.780.39 | 60.410.54 | 55.830.14 | |
| Local | 41.060.38 | 71.870.16 | 59.640.11 | 52.230.35 | 67.340.39 | 58.260.28 | |
| FL | 44.130.92 | 70.960.99 | 60.310.18 | 53.740.83 | 64.001.06 | 57.830.47 | |
| Local | 45.250.33 | 69.850.68 | 60.090.31 | 53.020.42 | 65.670.57 | 58.070.34 | |
| FL | 43.021.67 | 72.062.21 | 60.540.67 | 54.460.55 | 64.600.71 | 58.500.05 | |
| Local | 46.930.07 | 70.960.13 | 61.420.10 | 56.360.26 | 63.520.41 | 59.220.28 | |
| FL | 46.641.39 | 73.901.10 | 63.080.11 | 52.430.75 | 64.002.63 | 57.050.59 | |